OpenStudy (anonymous):

Given the system below: f(x)=2x g(x)=2x Which value(s) of x make f(x)=g(x) a true statement? If necessary, you many choose more than one answer. 0 1 2 3 4

OpenStudy (danjs):

f(x) = g(x) 2x = 2x what values can you use to keep that true,

OpenStudy (danjs):

you have two forms of the same line y=2x, they overlap at infinite points, and have infinite solutions

OpenStudy (chealyn98):

1) If f(x) = 2x and g(x) = 2x, then f(x) = g(x), since 2x = 2x. Therefore, f(x) = g(x) for all values of x, so all choices are correct (0, 1, 2, 3, 4). 2) For this one, you want to solve for y for both equations (the first one is already done): y = -x^2 - x + 2 x + y = -2 y = -x - 2 Now you can set both equations equal to each other: y = y -x^2 - x + 2 = -x - 2 -x^2 - x + x + 2 + 2 = 0 -x^2 + 4 = 0 -(x^2 - 4) = 0 x = -2, +2 To find the y-values, just plug x into either of the equations: y = -x - 2 y = -(-2) - 2 y = 2 - 2 y = 0 when x = -2 y = -2 - 2 y = -4 when x = +2 Your answers are (-2, 0) and (2, -4) 3) Solve this one using the exact same method as above: y = (1/2)x^2 + 2x - 1 3x - y = 1 -y = -3x + 1 y = 3x - 1 (1/2)x^2 + 2x - 1 = 3x - 1 (1/2)x^2 + 2x - 3x -1 + 1 = 0 (1/2)x^2 - x = 0 x(1/2x - 1) = 0 Now set each factor equal to 0 and solve for x: x = 0 (1/2)x - 1 = 0 (1/2)x = 1 x = 2 Plug x into either equation to get y: y = 3(0) - 1 y = -1 when x = 0 y = 3(2) - 1 y = 6 - 1 y = 5 when x = 2 The answers are (0, -1) and (2, 5)

OpenStudy (anonymous):

Ok thanks guys

OpenStudy (chealyn98):

Your welcome